A juice company sells cases of juice to sports camps at an average price of $13.62 and the materials cost an average of $8.48 per case. If they produce 7123 cases per week, what will their profit be if they sell 4489 cases during a particular week? (Unsold cases are thrown away or donated to a shelter, both of which result in $0 salvage value)
Answer #1

Profit = Revenue Earned - Cost price

Revenue Earned = 4489*13.62

= $61140.18

Cost Price = 7123*8.48

= $60403.04

Profit = 61140.18 - 60403.04

= $737.14